An aluminum plate is made according to the following dimensions. The outer radius is 10 inches and the inner radius is 3 inches. Use p = 3.14 as an approximate value. (a) How many square inches in area is the aluminum plate (b) If the cost to construct the plate is $7.50 per square inch, what is the construction cost?

Answer by nerdybill(5404)

An aluminum plate is made according to the following dimensions. The outer radius is 10 inches and the inner radius is 3 inches. Use p = 3.14 as an approximate value. (a) How many square inches in area is the aluminum plate (b) If the cost to construct the plate is $7.50 per square inch, what is the construction cost?
.
The aluminum plate is "round" (10 in radius) with a hole in the middle (hole has a 3 inch radius).
.
Part a:
"area of plate" = "area of circle w/10 in. radius" - "area of hole in middle"
.
"area of a circle" = (1/2)(pi)(r^2)
.
"area of plate" = [(1/2)(3.14)(10^2)] - [(1/2)(3.14)(3^2)]
"area of plate" = [(1/2)(3.14)(100)] - [(1/2)(3.14)(9)]
"area of plate" = (1/2)(3.14)[100-9]
"area of plate" = (1/2)(3.14)[91]
"area of plate" = 142.87 sq inches
.
Part b:
"cost" = 7.50(142.87) = $1071.53